OpenStudy (anonymous):

Simply the expression. Write your answer in lowest terms and leave it in factored form 2y/y(3y-1) + 1/ 3y-1

OpenStudy (anonymous):

to make the denominators the same, mutliply \(\frac{1}{3y-1}\) top and bottom by \(\frac{y}{y}\) to get \(\frac{y}{y(3y-1)}\) now that the denominators are the same, add up in the numerator

OpenStudy (anonymous):

So then it's just 2y+1= 3y/y(3y-1)

OpenStudy (anonymous):

the "final answer" is \(\frac{3y}{y(3y-1)}=\frac{3}{3y-1}\)